# Univariate-Guided Sparse Regression

### v8

@@ 3 Relationship of UniLasso to the Adaptive Lasso
Finally we note that ? propose an iterated version of the adaptive Lasso, in which the current solutions are used to define weights for the next iteration. They show that the method can enhance sparsity, especially for signal recovery.
4 Application to the diabetes data
Used in ?, this dataset consists of ten baseline variables: age, sex, body mass index, average blood pressure and six blood serum measurements. The measurements were obtained for each of $n=442$ diabetes patients, as well as the response of interest, a quantitative measure of disease progression one year after baseline.
Table 3 shows the full least squares fit on the left, and the Lasso and uniLasso coefficients in the middle panel. The uniLasso coefficients for both the unregularized case (UniReg, ($\lambda=0$) and the cv choice of $\lambda$ are shown; they are very similar. We did 200 half-samplings of the data, training on the first half and computing test error on the second half. The two rightmost columns of Table 3 shows the frequency of selection over the 200 realizations.
Least Squares
Estimate Std. Error t value Pr($>$||||t||||) Lasso UniReg UniLasso Lasso-freq UniLasso-freq
age -0.01 0.04 -0.17 0.87 0.00 0.00 0.00 124.00 4.00
sex -0.15 0.04 -3.92 0.00 -0.12 0.00 -0.00 196.00 4.00
bmi 0.32 0.04 7.81 0.00 0.32 0.34 0.33 200.00 200.00
bp 0.20 0.04 4.96 0.00 0.18 0.16 0.16 200.00 197.00
s1 -0.49 0.26 -1.90 0.06 -0.06 0.00 0.00 152.00 0.00
s2 0.29 0.21 1.41 0.16 0.00 0.00 0.00 60.00 0.00
s3 0.06 0.13 0.47 0.64 -0.14 -0.12 -0.10 184.00 191.00
s4 0.11 0.10 1.10 0.27 0.00 0.00 0.00 82.00 15.00
s5 0.46 0.11 4.37 0.00 0.32 0.30 0.33 200.00 200.00
s6 0.04 0.04 1.02 0.31 0.03 0.01 0.01 163.00 96.00
Table 3: Least squares, Lasso and UniLasso results on full data
It is interesting to first compare the unregularized models least squares and UniReg. The biggest surprise is the sex coefficient: it is negative and highly significant in least squares but is not even chosen by uniLasso! It turns out that sex on its own has a positive coefficient and is highly significant, and this explains why uniLasso did not choose it. Now what causes the sign change for sex? Due to the correlation between bmi and sex, adding sex to the model containing bmi has a strong effect. In addition, there is a moderately strong interaction between bmi and sex, with a negative coefficient. Hence the strong negative coefficient of sex in the LS model is really reflecting the relationship between bmi and sex. In summary, looking at both LS and unregularized uniLasso can help us to understand some hidden effects.
Turning now to lasso and unLasso with regulariization, Table 4 shows the test error: Least Squares (LS) has error about 1% lower than Lasso, and 6.5% lower than UniLasso. However as seen in Table 4, UniLasso displays much smaller average model size. Further in Table 5 we see that age appears in about 60% of the models chosen by LS and the lasso, and about half the time, its coefficient is negative. Features s2 and s3 show similar behavior.
LS Lasso UniLasso
mean 51.72 52.20 55.22
se 0.44 0.44 0.45
support size 10.00 7.80 4.54
Table 4: Average test error and support over 200 half-samples
Least Squares Lasso
Feature -1 0 1 -1 0 1
age+ 108 0 92 60 76 64
s2+ 17 0 183 28 140 32
$\bm{-}$ 58 0 142 158 16 26
Table 5: Three features with the most varying coefficient signs over the 200 half-samples. Their univariate signs are $+$,$+$,-- respectively.
Figure 5 shows the distribution of coefficients over 200 half-samplings. Those for UniLasso are much sparser and always have the same sign. In this example, the choice between Lasso and UniLasso depends on whether prediction error or model stability and interpretation are more important.
Figure 5: Coefficients of diabetes predictors over 200 half-samplings.
5 Analysis of UniLasso with Orthonormal Columns
@@
In this section, in the special case of orthonormal features, we derive explicit formula for the UniLasso coefficients. Figure 6 shows the Lasso and UniLasso paths for a simulated example with an orthonormal feature matrix. They look quite different, with the UniLasso path being considerable sparser.
Figure 6: Coefficient paths for Lasso and UniLasso, in a simulated example with 10 orthonormal features.
@@
$\min_{\theta_{j}}\frac{1}{2}\|y-\sum_{\ell=1}^{p}\theta_{\ell}{\hat{\beta}_{\ell}}x_{\ell}\|_{2}^{2}+\lambda\|\theta\|_{1}$
@@
Since the $x_{j}$ are orthogonal and unit norm, and using $x_{j}^{\top}y=\hat{\beta}_{j}$, we get ${\hat{\beta}_{j}}^{2}\theta_{j}={\hat{\beta}_{j}^{2}}-\lambda\cdot\text{sign}(\theta_{j}),$ and hence
@@
The SparseNet procedure (?) uses a thresholding function that has a roughly similar shape to that of UniLasso, but its objective is not convex. This shrinkage pattern is somewhere between Lasso and $\ell_{0}$ or best subset selection. UniLasso achieves this without losing convexity, a significant computational advantage.
@@ 6 Theoretical analysis
$\hat{\beta}_{i,j}:=\frac{\frac{1}{n-1}\sum_{k\neq i}Y_{k}X_{k,j}-(\frac{1}{n-1}\sum_{k\neq i}Y_{k})(\frac{1}{n-1}\sum_{k\neq i}X_{k,j})}{\frac{1}{n-1}\sum_{k\neq i}X_{k,j}^{2}-(\frac{1}{n-1}\sum_{k\neq i}X_{k,j})^{2}}$
@@
$\hat{\alpha}_{i,j}:=\frac{1}{n-1}\sum_{k\neq i}Y_{k}-\frac{\hat{\beta}_{i,j}}{n-1}\sum_{k\neq i}X_{k,j}$
be the intercept term. Note that $\hat{\beta}_{i,j}$ is a consistent estimate of
@@
and $\hat{\alpha}_{i,j}$ is a consistent estimate of
@@
The above theorem is roughly comparable to the available results for the Lasso. A close comparison would be, for instance, (?, Theorem 11.3). Like Theorem 1, this theorem also assumes a lower bound on the covariance matrix of the covariates in the support, and the maximum difference between $\hat{\gamma}_{j}$ and $\gamma_{j}$ for $j\in S$ is of order $\lambda$. However, there is one key difference. The results about Lasso, including the one cited above, require a condition known as mutual incoherence or irrepresentability. Roughly speaking, it means that if we regress $X_{k}$ for some $k\notin S$ on $(X_{j})_{j\in S}$, the regression coefficients should be small. Notably, our Theorem 1 requires no such relation to hold between the covariates inside and outside the support. All we need is that the univariate regression coefficients of $Y$ on covariates outside the support are small.
@@ 7 Some simulation results
Figures 8 shows the results of a comparative study of UniLasso with Lasso and two other sparse regression methods. Shown are means and standard errors over 100 simulations.
The methods are:
  • Lasso (using glmnet),
@@
  • Adaptive Lasso using penalty weights $1/|\hat{\beta}_{j}^{uni}|$. These first three methods use CV to estimate the value of $\lambda$ that minimizes test error.
@@
1. Low, medium and high SNR: Here $n=300,p=1000$, $N(0,1)$ features with pairwise correlation 0.5; The nonzero regression coefficients are $N(0,1)$; Gaussian errors with SD $\sigma$ chosen to produce low $<1$, medium ($\approx 1$) , or high (>2)>2)> 2 ) SNRs.
2. Homecourt: As in #1, but the feature covariance is AR(1) with $\rho=0.8$ and the non-zero regression coefficients are $U(0.5,2)$, and are positioned on every other feature $(1,3,\ldots 99)$.
3. Two-class: The feature covariance is in #2, with n = 200, p = 500 and a binary target y. The first 20 features are shifted in the y = 1 class by 0.5 units.
@@
In the first 3 scenarios, 10% of the true coefficients are non-zero. Figure 8 shows the test set error (relative to the Lasso) and support size (number of non-zero coefficients)
Figure 8: Blue lines: relative test error (Gaussian) , misclassification rate (Binomial) rate or deviance (Cox), with support size superimposed in gold.
Overall we see that UniLasso shows MSE similar to that of Lasso, with often a much smaller support. The same general pattern holds with the real datasets.
Exceptions to this are the high-SNR setting where Lasso wins handily and the homecourt setting designed to exploit UniLasso’s strength (recall Theorem 2). In neutral court, the regression coefficients are allowed to have arbitrary signs and the advantage of UniLasso goes away. The “counter-example” setting is the classic case where features have positive correlation but the true regression coefficients have opposite signs; not surprisingly UniLasso does poorly. Neither the adaptive Lasso or the matching method show much advantage, compared to Lasso and UniLasso .
@@
Figure 9 shows the corresponding results for the simulated examples, for $n=300,p=100$.
Figure 9: $N>p$ case. Blue lines: relative test error (Gaussian) misclassification rate (Binomial) or deviance (Cox), with support size superimposed in gold.
@@ 8.1 Simulated data
Here we consider the case where $\lambda=0$, so that there is no $\ell_{1}$ penalty, as detailed in equation (4). While the main use case has $n>p$, we note that uniReg can be defined for $p>n$ by taking $\lambda\rightarrow 0$ in uniLasso. Computationally, this will approximately give the solution with minimum $\ell_{1}$ norm, in the same way that glmnet gives the least squares solution with minimum $\ell_{1}$ norm.
Conveniently even without the $\ell_{1}$ penalty, the non-negativity constraint still promotes sparsity in the solution. In Table 6 we compare least squares with uniReg in a subset of the simulated settings defined in Section 7. UniReg performs remarkably well, often winning in MSE and yielding consistently sparser models.
@@
$N=300,p=100$
MSE
setting family n p SNR mse-LS mse-uniReg
1 low-SNR gaussian 300 100 0.318 97.678 69.655
2 medium-SNR gaussian 300 100 1.272 24.428 18.559
3 high-SNR gaussian 300 100 5.087 6.104 5.82
4 homecourt gaussian 300 100 0.228 149.683 104.979
Support
setting family n p SNR supp-LS supp-uniReg
1 low-SNR gaussian 300 100 0.318 100 8
2 medium-SNR gaussian 300 100 1.272 100 9
3 high-SNR gaussian 300 100 5.087 100 7
4 homecourt gaussian 300 100 0.228 100 13
$N=300,p=30$
MSE
setting family n p SNR mse-LS mse-uniReg
1 low-SNR gaussian 300 30 0.092 71.339 65.171
2 medium-SNR gaussian 300 30 0.369 17.836 16.489
3 high-SNR gaussian 300 30 1.475 4.459 4.178
4 homecourt gaussian 300 30 0.059 111.771 101.542
Support
setting family n p SNR supp-LS supp-uniReg
1 low-SNR gaussian 300 30 0.092 30 4
2 medium-SNR gaussian 300 30 0.369 30 4
3 high-SNR gaussian 300 30 1.475 30 4
4 homecourt gaussian 300 30 0.059 30 4
$N=300,p=1000$
MSE
setting family n p SNR mse-LS mse-uniReg
1 low-SNR gaussian 300 1000 0.578 842.191 510.837
2 medium-SNR gaussian 300 1000 1.48 374.915 249.076
3 high-SNR gaussian 300 1000 3.613 186.171 147.276
4 homecourt gaussian 300 1000 1.53 577.852 420.508
Support
setting family n p SNR supp-LS supp-uniReg
1 low-SNR gaussian 300 1000 0.578 1000 24.5
2 medium-SNR gaussian 300 1000 1.48 1000 27
3 high-SNR gaussian 300 1000 3.613 1000 27
4 homecourt gaussian 300 1000 1.53 1000 135.5

### v9

@@ 3 Relationship of uniLasso to the adaptive lasso
Finally we note that Candes et al. (2008) propose an iterated version of the adaptive lasso, in which the current solutions are used to define weights for the next iteration. They show that the method can enhance sparsity, especially for signal recovery.
Next we examine the relative performance of uniLasso when we modify the three main aspects of its design:
  1. (a)
    the use of LOO features $\hat{\eta}_{j}^{-i}$ versus non-LOO features $\hat{\eta}_{j}^{i}$,
  2. (b)
    the sign constraints, and
  3. (c)
    the scaling in Step 2 by the magnitude of the univariate coefficients, rather than just by their signs.
Note that the adaptive lasso is equivalent to uniLasso using non-LOO estimates and no sign constraints.
Table 3 shows the results of 50 simulation runs with $n=300,\;p=1000$ and SNR=1, our “medium-SNR” setting detailed in Section 8.
LOO
lasso uniLasso uniLasso-noSign uniLasso-noMag
MSE 0.55 0.59 0.61 0.60
se 0.03 0.03 0.04 0.04
Support 53.78 15.32 66.38 40.12
se 4.41 1.19 12.92 6.47
No LOO
lasso uniLasso uniLasso-noSign uniLasso-noMag
MSE 0.55 0.61 0.57 0.61
se 0.03 0.04 0.03 0.05
Support 53.78 36.84 24.48 48.98
se 4.41 6.34 2.71 9.66
Table 3: Results for 50 simulations from the medium-SNR setting. The last two columns use variations of uniLasso: noSign removes the non-negative constraint, while noMag uses just the sign of the univariate estimates in Step 2 of uniLasso (not the magnitude). The top table uses LOO univariate coefficients (as in uniLasso) while the bottom table uses the usual (non-LOO) versions.
We see that there is not much difference in test error, but uniLasso produces the sparsest models by a large margin. We also see the curious result that with no LOO, the uniLasso without sign constraints selects sparser models than with sign constraints.
4 Why does the uniLasso produce such sparse solutions?
As we saw in the previous section, the combination of LOO univariate estimates in Step (1) and nonnegative constraints in step(2) seem to be the key to the strong sparsity delivered by uniLasso. What is the explanation for this?  [^2] We thank Chris Habron for this analysis.
The correlation between the responses $y_{i}$ and the univariate fitted values $\hat{\beta}_{0j}+\hat{\beta}_{j}x_{ij}$ is always positive, and is the absolute value of the correlation between $y$ and the $j$th feature. When this correlation is large, then the correlation between $y_{i}$ and the LOO fitted values $\hat{\beta}_{0j}^{-i}+\hat{\beta}_{j}^{-i}x_{ij}$ will also tend to be positive. But if the first correlation is positive and small then the second correlation is often negative. This could be explained by the fact that leaving an observation out “pushes” the fitted regression line away from that point, and if there wasn’t much correlation to start with that will be enough to tip the correlation of the LOO feature to be negative. [^3] In the extreme case of zero correlation between $y$ and the $j$th feature, the univariate fit is $\hat{\beta}_{0j}=\bar{y}$, and one can show that the the correlation between $y$ and the LOO version of $\bar{y}$ is -1!
Figure 2 shows an example from our “medium-SNR” setting with $n=300,p=1000$. The correlations between $y$ and each feature needs to be larger than about 0.08 in order for the correlation between $y$ and each LOO feature to also be positive.
Figure 2: Medium-SNR setting: correlation between the response and the univariate LOO features, versus the correlation between the response and the usual univariate (non-LOO) features. The solid line represents equality.
In Appendix D we show that if the absolute correlation between $y$ and a feature is smaller than approximately $\sqrt{2/n}$, then the correlation with the LOO fit will be negative. Since we have $n=300$ in this example, this evaluates to $0.082$. Asymptotically, this means that the two-sided $p$-value for testing $\rho=0$ has to greater than about 0.16.
In the second step of the uniLasso algorithm in Section 2.1, we fit a multivariate lasso model using all the LOO features, but with a positivity constraint on their coefficients. So even though marginally some LOO features are negatively correlated with the response, this does not guarantee that they will be omitted; however, in practice they do tend to be omitted.
5 Application to the car prices data
This dataset consists of 25 features for predicting car prices for 205 cars, taken from https://www.kaggle.com/datasets/hellbuoy/car-price-prediction Make of car (22 levels) was first fit to the data and we modeled the residual on the remaining 26 predictors, 11 of which were categorical (and one-hot encoded).
Univariate Lasso UniLasso Polish
body.style-3 -4.768 -0.224 0.000 0.000
wheel.base 0.781 0.035 0.000 0.000
width 3.082 0.475 0.378 0.000
height 0.326 -0.154 0.000 0.000
curb.weight 0.014 0.004 0.002 -0.001
engine.type-7 -0.652 3.303 0.000 0.000
num.of.cylinders-2 9.867 -2.546 0.000 0.000
num.of.cylinders-3 -13.854 $0$0.545 0.000
num.of.cylinders-4 10.543 -0.667 0.000 0.000
num.of.cylinders-7 -0.652 0.009 0.000 0.000
engine.size 0.167 0.064 0.048 0.000
fuel.system-2 -9.143 0.208 0.000 0.000
fuel.system-5 -0.694 -1.248 0.000 0.000
bore 17.571 -3.729 0.000 0.000
horsepower 0.169 0.017 0.042 0.000
peak.rpm -0.001 0.000 0.000 0.001
Table 4: Coefficients from various models fit to the car-price data. Coefficients with sign changes relative to the univariate fits are marked in blue.
Table 4 shows the univariate least squares coefficients on the left, and the lasso and uniLasso coefficients in the middle columns. The right column shows the “uniLasso polish” estimates described in Section 13. They result from a post-processing of uniLasso in which the lasso is applied to the uniLasso residuals. Only features having a non-zero coefficient in at least one of the three rightmost columns are shown. We see that uniLasso produces a much sparser model than the lasso.
For Figure 3 we took 50 random $2/3-1/3$ train-test splits, and computed 4 summaries of model performance. We see that the test errors of all methods are similar, while the support size and number of sign-violations are quite different, as expected. The stability figure is most interesting: for each of the (${50\choose 2}$ pairs of models, we defined the “stability” as ratio of the number of chosen features in common, divided by the number of unique features in the union of the pair. We see a clear advantage for uniLasso over lasso, where the median proportion of common features is over 80%, versus 60% for lasso. The uniLasso “polish”, described in Section 13 gives a further improvement in stability, while increasing the support and number of sign violations.
Figure 3: Results for car prices data over 50 train-test splits. Shown are test set error, size of the chosen model (support), number of sign change violations relative to the univariate signs, and the stability — the average proportion of common features over all model pairs for each method.
6 Analysis of uniLasso with orthogonal features
@@
In this section we derive explicit formula for the uniLasso coefficients in the special case of orthonormal features. Figure 4 shows the lasso and uniLasso paths for a simulated example with an orthonormal feature matrix. They look somewhat different, with the uniLasso path being sparser at any stage along the path (as measured by the $\ell_{1}$ norm of the coefficients).
Figure 4: Coefficient paths for lasso and uniLasso, in a simulated example with 10 orthonormal features. The end of the path represents the least squares fit, and in this case also the univariate coefficients. The paths for features with large absolute univariate coefficients look very similar in both plots. Those with small univariate features have a delayed entry in the right plot.
@@
$\min_{\theta_{j}}\left\{\frac{1}{2}\|y-\sum_{\ell=1}^{p}\theta_{\ell}{\hat{\beta}_{\ell}}x_{\ell}\|_{2}^{2}+\lambda\|\theta\|_{1}\right\}$
@@
Since the $x_{j}$ are orthogonal and unit norm, and using $x_{j}^{\top}y=\hat{\beta}_{j}$, we get
${\hat{\beta}_{j}}^{2}\theta_{j}={\hat{\beta}_{j}^{2}}-\lambda\cdot\text{sign}(\theta_{j}),$ and hence
@@
The SparseNet procedure (Mazumder et al., 2011) uses a thresholding function that has a roughly similar shape to that of uniLasso, but its objective is not convex. This shrinkage pattern is somewhere between lasso and $\ell_{0}$ or best subset selection. UniLasso achieves this without losing convexity, a significant computational advantage.
@@ 7 Theoretical analysis
$\hat{\beta}^{-i}_{j}:=\frac{\frac{1}{n-1}\sum_{k\neq i}Y_{k}X_{k,j}-(\frac{1}{n-1}\sum_{k\neq i}Y_{k})(\frac{1}{n-1}\sum_{k\neq i}X_{k,j})}{\frac{1}{n-1}\sum_{k\neq i}X_{k,j}^{2}-(\frac{1}{n-1}\sum_{k\neq i}X_{k,j})^{2}}$
@@
$\hat{\alpha}^{-i}_{j}:=\frac{1}{n-1}\sum_{k\neq i}Y_{k}-\frac{\hat{\beta}^{-i}_{j}}{n-1}\sum_{k\neq i}X_{k,j}$
be the intercept term. Note that $\hat{\beta}^{-i}_{j}$ is a consistent estimate of
@@
and $\hat{\alpha}^{-i}_{j}$ is a consistent estimate of
@@
The above theorem is roughly comparable to the available results for the lasso. A close comparison would be, for instance, (Hastie et al., 2015, Theorem 11.3). Like Theorem 1, this theorem also assumes a lower bound on the covariance matrix of the covariates in the support, and the maximum difference between $\hat{\gamma}_{j}$ and $\gamma_{j}$ for $j\in S$ is of order $\lambda$. However, there is one key difference. The results about lasso, including the one cited above, require a condition known as mutual incoherence or irrepresentability. Roughly speaking, it means that if we regress $X_{k}$ for some $k\notin S$ on $(X_{j})_{j\in S}$, the regression coefficients should be small. Notably, our Theorem 1 requires no such relation to hold between the covariates inside and outside the support. All we need is that the univariate regression coefficients of $Y$ on covariates outside the support are small.
@@ 8 Some simulation results
Figure 6 shows the results of a comparative study of uniLasso with lasso and three other sparse regression methods. Shown are means and standard errors over 100 simulations. The methods are:
  • Lasso (using glmnet).
@@
  • Polish: a post-processing of the uniLasso solution, described in Section 13.
  • Adaptive lasso using penalty weights $1/|\hat{\beta}_{j}|$ using the univariate coefficients $\hat{\beta}_{j}$. These first three methods use CV to estimate the value of $\lambda$ that minimizes test error.
@@
1. Low, medium and high SNR: Here $n=300$, $p=1000$, $N(0,1)$ features with pairwise correlation 0.5. The nonzero regression coefficients are $N(0,1)$; Gaussian errors with SD $\sigma$ chosen to produce low $<1$, medium ($\approx 1$) , or high (>2)>2)> 2 ) SNRs.
2. Homecourt: Here $n=100$, $p=30$ with details as specified in Section 2.3.
3. Two-class: We set $n=200$, $p=500$ with a binary target $y$. The feature covariance within each class is AR(1) with $\rho=0.8$, and the first 20 features are shifted in the $y=1$ class by 0.5 units.
@@
In the first three scenarios, 10% of the true coefficients are non-zero; for example, in (1), 100 of the 1000 true coefficients are non-zero. Figure 6 shows the test set error (relative to the lasso) and support size (number of non-zero coefficients).
Figure 6: $N=300$, $p=1000$: test error (Gaussian) or misclassification rate (Binomial), along with support size.
Overall we see that uniLasso shows MSE similar to that of lasso, with often a much smaller support. The same general pattern holds with the real datasets. Exceptions to this are the high-SNR setting where lasso wins handily and the homecourt setting designed to exploit uniLasso’s strength (recall Theorem 2). The “counter-example” setting is the classic case where features have positive correlation but the true regression coefficients have opposite signs; not surprisingly uniLasso does poorly. The polish method and adaptive lasso do reasonably well, while the matching method does poorly.
@@
Figure 7 shows the corresponding results for the simulated examples with $N=300$ and $p=100$.
Figure 7: $N=300$$p=100$: Test error (Gaussian) or misclassification rate (Binomial), along with support size.
@@ 9 The unregularized setting: uniReg
Here we consider the case where $\lambda=0$, so that there is no $\ell_{1}$ penalty, as detailed in equation (1). While the main use case has $n>p$, we note that uniReg can be defined for $p>n$ by taking $\lambda\rightarrow 0$ in uniLasso. Computationally, this will approximately give the solution with minimum $\ell_{1}$ norm, in the same way that the limiting lasso fit gives the least squares solution with minimum $\ell_{1}$ norm. Conveniently even without the $\ell_{1}$ penalty, the non-negativity constraint still promotes sparsity in the solution.
In this section we demonstrate the performance of uniReg on simulated and real data.
@@ 9.1 Simulated data
In Table 5 we compare least squares with uniReg in a subset of the simulated settings defined in Section 8. UniReg performs remarkably well, often winning in MSE and yielding consistently sparser models.
N=300, p=30
Setting p SNR MSE-LS MSE-uniReg Supp-LS Supp-uniReg
low-SNR 30 0.092 71.339 65.171 30 4
medium-SNR 30 0.369 17.836 16.489 30 4
high-SNR 30 1.475 4.459 4.178 30 4
homecourt 30 0.059 111.771 101.542 30 4
N=300, p=100
Setting p SNR MSE-LS MSE-uniReg Supp-LS Supp-uniReg
low-SNR 100 0.318 97.678 69.655 100 8
medium-SNR 100 1.272 24.428 18.559 100 9
high-SNR 100 5.087 6.104 5.82 100 7
homecourt 100 0.228 149.683 104.979 100 13
N=300, p=1000
Setting p SNR MSE-LS MSE-uniReg Supp-LS Supp-uniReg
low-SNR 1000 0.578 842.191 510.837 300 24.5
medium-SNR 1000 1.48 374.915 249.076 300 27
high-SNR 1000 3.613 186.171 147.276 300 27
homecourt 1000 1.53 577.852 420.508 300 135.5